# ChainMarks: Securing DNN Watermark with Cryptographic Chain

### v1

@@ Abstract.
conference: ACM ASIA Conference on Computer and Communications Security; 25-29 August, 2025; Ha Noi, Vietnamjournalyear: 2025copyright: rightsretainedconference: ACM Asia Conference on Computer and Communications Security; August 25–29, 2025; Ha Noi, Vietnambooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’25), August 25–29, 2025, Ha Noi, Vietnamdoi: 10.1145/xxxxxxx.xxxxxxxisbn: xxx-x-xxxx-xxxx-x/xx/xxccs: Security and privacy Security services
@@ 10.3. Attacks/Defenses on DNN Watermarks
Model Modification Attacks. After a model is pirated, the model weights are often modified by the pirate. Model modification attacks include model fine-tuning (Uchida et al., 2017; Chen et al., 2019c, 2021b), model pruning or parameter pruning (Rouhani et al., 2019), model weight compression (Uchida et al., 2017), and model retraining (Cao et al., 2021; Namba and Sakuma, 2019).
Evasion Attacks. Evasion attacks are more complicated than model modification attacks. Shafieinejad et al. investigate watermark removal attacks against backdoor-based watermarks using white-box, black-box, and inference attacks (Shafieinejad et al., 1906). A DNN laundering scheme can be leveraged to reset the backdoor watermarks (Aiken et al., 2021). Attackers can also manipulate the model to remove the owner’s signature if they know about the presence of the watermark in advance (Guo and Potkonjak, 2018). Reverse engineering may be used if attackers can obtain the original training dataset (Fan et al., 2019). Liu et al. propose a data augmentation scheme to imitate the behavior of the backdoor triggers (Liu et al., 2021). Gong et al. propose a dynamic learning rate adjustment strategy to purify the backdoors (Gong et al., 2023). Attackers can leverage resource-efficient attack techniques (Ilyas et al., 2018; Tu et al., 2019; Suya et al., 2020) to find black-box adversarial examples that are even transferable between different DNN models. By combining hybrid attacks with the seed prioritization strategy, adversarial examples can be obtained using only a few queries. The two main optimization techniques used in attacks are AutoZOOM (Tu et al., 2019) and NES (Ilyas et al., 2018).

### v2

@@ Abstract.
conference: ACM ASIA Conference on Computer and Communications Security; 25-29 August, 2025; Ha Noi, Vietnamjournalyear: 2025copyright: ccconference: ACM Asia Conference on Computer and Communications Security; August 25–29, 2025; Hanoi, Vietnambooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’25), August 25–29, 2025, Hanoi, Vietnamdoi: 10.1145/3708821.3736214isbn: 979-8-4007-1410-8/2025/08ccs: Security and privacy Security services
@@ 10.3. Attacks/Defenses on DNN Watermarks
Model Modification Attacks. Model weights are often modified by the pirate. Model modification attacks include model fine-tuning (Uchida et al., 2017; Chen et al., 2019c, 2021b), model pruning or parameter pruning (Rouhani et al., 2019), model weight compression (Uchida et al., 2017), and model retraining (Cao et al., 2021; Namba and Sakuma, 2019).
Evasion Attacks. Evasion attacks are more complicated. Shafieinejad et al. investigate the removal of backdoor-based watermarks with white-box, black-box, and inference attacks (Shafieinejad et al., 1906). Also, DNN laundering is shown to reset backdoor watermarks (Aiken et al., 2021). Attackers can manipulate a model to remove the owner’s signature if watermark presence is known in advance (Guo and Potkonjak, 2018). Reverse engineering can be used if the original training dataset is obtained (Fan et al., 2019). Liu et al. propose a data augmentation scheme to mimic the backdoor trigger behaviors (Liu et al., 2021). Gong et al. dynamically adjust the learning rate to purify backdoors (Gong et al., 2023). Attackers can leverage resource-efficient attacks (Ilyas et al., 2018; Tu et al., 2019; Suya et al., 2020) to find black-box adversarial examples. By combining hybrid attacks with seed prioritization, adversarial examples can be obtained using only a few queries. The two main optimization techniques used in attacks are AutoZOOM (Tu et al., 2019) and NES (Ilyas et al., 2018).
@@ 11. Conclusion
Acknowledgements.
This work is partially supported by Office of Naval Research grant N00014-23-1-2122 and Commonwealth Cyber Initiative award N-3Q25-002.